how much pure acid must be added to 25 ounces of a solution of acid and water that is 20% pure acid in order to make a solution that is 50 % pure acid ?
Answer by ewatrrr(24785)


Hi
Let x represent how much pure acid needs to added to the 20% solution
Question states***
"make a solution that is 50 % pure acid"
1.0x + .20*25oz = .50(25oz+x)
.5x = .30*25oz
x = 15 oz pure acid needs to added to the 20% solution
CHECKING our Answer***
1.0*15oz + .20*25oz = 15oz + 5oz = .5*40 = 20oz (pure acid)